OpenStudy (anonymous):

what is the domain if the range is all real numbers? f(x) = the square root of 6-x

OpenStudy (anonymous):

\[\sqrt{6-x}\]

jimthompson5910 (jim_thompson5910):

close but no

jimthompson5910 (jim_thompson5910):

you claim the domain is x <= -6, but I can plug in x = 0 y = sqrt(6 - x) y = sqrt(6 - 0) y = sqrt(6) and I get a real number just fine. So that's not the whole domain.

jimthompson5910 (jim_thompson5910):

you need to solve 6-x >= 0 for x

jimthompson5910 (jim_thompson5910):

it sounds like you're guessing at this point

jimthompson5910 (jim_thompson5910):

solve 6-x >= 0 for x if you multiply both sides of the inequality by a negative number, then the inequality sign will flip if you divide both sides of the inequality by a negative number, then the inequality sign will flip

OpenStudy (anonymous):

so thats x>=6

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

why not add x to both sides

jimthompson5910 (jim_thompson5910):

that may be easier

OpenStudy (anonymous):

6-x>=0 -6 -6 --------- -x>-6 -- ---- -1 -1 x<=6

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

yes I already said "correct" above lol